A 3.5kg block is pushed along a horizontal floor by a force F of magnitude 15N at an angle 40°W horizontal. The coefficient of kinetic friction between the block and the floor is 0.25. Calculate the magnitude of:



(a) frictional force on the block from the floor



(b) the block's acceleration

1
Expert's answer
2022-03-10T10:17:41-0500


The Newton's second law says

"ma=F\cos40^\circ-F_f"

and

"N=mg-F\sin40^\circ\\
=3.5*9.8-15\sin40^\circ=24.7\:\rm N"

Hence:

(a) frictional force on the block from the floor

"F_f=\mu N=0.25*24.7=6.2\:\rm N"

(b) the block's acceleration

"a=(15\cos40^\circ-6.2)/3.5=1.5\:\rm m/s^2"